Hook Industries's capital structure consists solely of debt and common equity. It can issue debt at rd = 11%, and its common stock currently pays a $3.75 dividend per share (D0 = $3.75). The stock's price is currently $33.50, its dividend is expected to grow at a constant rate of 6% per year, its tax rate is 35%, and its WACC is 13.30%. What percentage of the company's capital structure consists of debt? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
33.5 = 3.75 * (1+0.06) / (Cost of equity - 0.06)
Cost of equity% = 17.87
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 11*(1-0.35)
= 7.15
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
13.3=7.15*(weight of debt)+17.87*(1-weight of debt)

Weight of debt = 42.63%